# 1st Arshia Hemmat dept. Computer Engineering

### v1

@@ Abstract
This paper introduces an innovative approach using Retrieval-Augmented Generation (RAG) pipelines with Large Language Models (LLMs) to enhance information retrieval and query response systems for university-related question answering. By systematically extracting data from the university’s official webpage and employing advanced prompt engineering techniques, we generate accurate, contextually relevant responses to user queries.
We developed a comprehensive university benchmark, UniversityQuestionBench (UQB), to rigorously evaluate our system’s performance, based on common key metrics in the filed of RAG pipelines, assessing accuracy and reliability through various metrics and real-world scenarios. Our experimental results demonstrate significant improvements in the precision and relevance of generated responses, enhancing user experience and reducing the time required to obtain relevant answers. In summary, this paper presents a novel application of RAG pipelines and LLMs, supported by a meticulously prepared university benchmark, offering valuable insights into advanced AI techniques for academic data retrieval and setting the stage for future research in this domain.
@@ II-B Recent Advances and Techniques in RAG
Recent advancements in RAG have focused on innovative techniques and methodologies to optimize retrieval and generation processes. Lewis et al. (2020) highlight the power of RAG in knowledge-intensive NLP tasks, demonstrating its potential to solve complex information retrieval challenges [1]. Shahul et al. (2023) introduced RAGAS, a framework for automated evaluation of RAG pipelines, emphasizing the importance of reference-free evaluation metrics to enhance the evaluation process of RAG systems. Siriwardhana et al. (2022) developed RAG-end2end, which optimizes RAG for domain-specific knowledge bases, significantly improving performance in specialized domains such as healthcare and news [4]. Yu (2022) explored the use of retrieval-augmented generation across heterogeneous knowledge, addressing the challenges of retrieving information from diverse sources [6]. Nakhod (2023) proposed applying RAG to elevate low-code developer skills by integrating domain-specific knowledge into large language models, thereby improving their practical utility [9]. Melz (2023) introduced ARM-RAG, a system that enhances large language models’ intelligence through storing and retrieving reasoning chains, demonstrating significant improvements in problem-solving tasks [10]. Chen et al. (2023) provided a comprehensive evaluation of the impact of RAG on large language models, highlighting the potential bottlenecks and challenges in applying RAG across different tasks [7].
@@ III-A Data Production Process
Step 2: Let $\mathcal{Q}_{\text{student}}$ represent the set of questions gathered from student surveys. The second step is formulated in the Equation 2:
@@
The dataset incorporates information from all departments at the University of Isfahan, ensuring comprehensive coverage of student queries. We illustrate the process of the Dataset Generation in the figure 2.
@@ VI-A Pipeline Performance On UQB
Our pipeline’s performance was evaluated on a dataset of 300 questions and answers using the test set of the UniversityQuestionBench (UQB) dataset. We used our base model, Dorna, to compute the evaluation metrics. As previously discussed, the evaluation focused on three key metrics: faithfulness, answer relevance, and context relevance. The calculated results are demonstrated in table I :
@@ VI-B Quality of Outputs
To assess the acceptability of the pipeline outputs from a human perspective, we conducted a qualitative evaluation involving subject matter experts and university students. The evaluators rated the generated answers based on clarity, coherence, and overall satisfaction. The feedback indicated that the majority of the answers were clear, well-structured, and effectively addressed the questions, demonstrating high acceptability.
Additionally, the consistency of the answers was noted, with evaluators observing a uniform level of quality across different questions. This consistency underscores the robustness of our pipeline, confirming that it produces reliable and high-quality answers suitable for practical use in a university setting.

### v2

@@ Abstract
This paper introduces an innovative approach using Retrieval-Augmented Generation (RAG) pipelines with Large Language Models (LLMs) to enhance information retrieval and query response systems for university-related question answering. By systematically extracting data from the university’s official website, primarily in Persian, and employing advanced prompt engineering techniques, we generate accurate and contextually relevant responses to user queries.
We developed a comprehensive university benchmark, UniversityQuestionBench (UQB), to rigorously evaluate our system’s performance. UQB focuses on Persian-language data, assessing accuracy and reliability through various metrics and real-world scenarios. Our experimental results demonstrate significant improvements in the precision and relevance of generated responses, enhancing user experiences, and reducing the time required to obtain relevant answers.
In summary, this paper presents a novel application of RAG pipelines and LLMs for Persian-language data retrieval, supported by a meticulously prepared university benchmark, offering valuable insights into advanced AI techniques for academic data retrieval and setting the stage for future research in this domain. [^1] Dataset is publicly available at https://huggingface.co/datasets/UIAIC/UQB
@@ II-B Recent Advances and Techniques in RAG
Recent advancements in RAG have focused on innovative techniques and methodologies to optimize retrieval and generation processes. Lewis et al. (2020) highlight the power of RAG in knowledge-intensive NLP tasks, demonstrating its potential to solve complex information retrieval challenges [1]. Shahul et al. (2023) introduced RAGAS, a framework for automated evaluation of RAG pipelines, emphasizing the importance of reference-free evaluation metrics to enhance the evaluation process of RAG systems. Siriwardhana et al. (2022) developed RAG-end2end, which optimizes RAG for domain-specific knowledge bases, significantly improving performance in specialized domains such as healthcare and news [4]. Yu (2022) explored the use of retrieval-augmented generation across heterogeneous knowledge, addressing the challenges of retrieving information from diverse sources [6]. Nakhod (2023) proposed applying RAG to elevate low-code developer skills by integrating domain-specific knowledge into large language models, thereby improving their practical utility [9]. Melz (2023) introduced ARM-RAG, a system that enhances large language models’ intelligence through storing and retrieving reasoning chains, demonstrating significant improvements in problem-solving tasks [10]. Chen et al. (2023) provided a comprehensive evaluation of the impact of RAG on large language models, highlighting the potential bottlenecks and challenges in applying RAG across different tasks [7]. Heydari et al. (2024) proposed the Context Awareness Gate (CAG) architecture, a novel mechanism that dynamically adjusts the LLM’s input prompt based on whether the user query necessitates external context retrieval, thereby enhancing the efficiency and accuracy of RAG systems [11].
@@ III-A Data Production Process
Step 2: In the second stage, we collected the questions from students through two primary approaches to ensure comprehensive coverage of frequently asked questions:
  1. 1.
    Student Surveys:We utilized data collected from 60 students to refine and enhance the quality of the documents, ensuring they accurately reflected the most relevant and practical information based on real student experiences and inquiries. We asked students to fill out a form regarding the most frequent questions they encounter at the university. The form included open-ended sections where students could describe the challenges they face when navigating university resources. From the responses, we identified a correspondence corpus of frequently asked questions that were representative of common student concerns from either the university website or the relevant documents in the different channels.
  2. 2.
    Web Scraping: In addition to the student surveys, we scraped the official Isfahan University website for relevant information. This included extracting questions derived from useful content such as department-specific pages, contact information (e.g., email addresses of professors), course descriptions, and administrative procedures. These questions were designed to complement the student-provided data and fill in any gaps in coverage.
The collected data from these two sources forms the initial set of questions, denoted as $\mathcal{Q}_{\text{student}}$. This step is formulated in Equation 2:
@@
The final dataset comprises over 500 documents, encompassing a wide range of contexts to ensure comprehensive coverage. These documents include detailed information about the academic groups and sections of each department, professors along with their contact emails, and various aspects of the university, such as administrative procedures, facilities, and other relevant resources. This extensive collection ensures that the dataset accurately represents the diverse informational needs of students across the university. We illustrate the process of the Dataset Generation in the figure 2.
@@ VI-A Pipeline Performance On UQB
Our pipeline’s performance was evaluated on a dataset of 300 questions and answers using the test set of the UniversityQuestionBench (UQB) dataset. We used our base model, Dorna, to compute the evaluation metrics. As previously discussed, the evaluation focused on three key metrics: faithfulness, answer relevance, and context relevance. The calculated results are demonstrated in Table I.
  • Faithfulness:
    • Faithfulness measures the factual accuracy of the responses, reflecting the system’s ability to generate outputs consistent with the underlying data source.
    • The highest faithfulness score (0.8497) is achieved by GPT-3.5-turbo with OpenAI Embeddings, underscoring the robustness of general-purpose embeddings in generating accurate responses.
    • The performance of Dorna with Dorna Embeddings (0.839) is competitive, highlighting the capability of localized embeddings specifically designed for Persian-language content.
    • Models utilizing Persian-Sentence-Embedding-V3 (e.g., GPT-3.5-turbo, 0.8113) exhibit slightly reduced faithfulness compared to OpenAI Embeddings, suggesting limitations in the current iteration of these embeddings for ensuring strict fidelity to source information.
  • Answer Relevancy:
    • Answer relevancy assesses the alignment of the generated response with the user’s query, a critical factor for user satisfaction.
    • Dorna with Dorna Embeddings outperforms other models with a score of 0.823, demonstrating its strength in producing highly relevant responses tailored to Persian-language queries. This indicates the effectiveness of the Dorna model in embedding semantic understanding of user intent in Persian.
    • The lowest score for this metric is observed with GPT-3.5-turbo using Persian-Sentence-Embedding-V3 (0.493), suggesting potential challenges in aligning query semantics with response generation when utilizing this embedding.
  • Context Relevancy:
    • Context relevancy measures how well the generated response incorporates broader contextual understanding, ensuring a coherent and comprehensive answer.
    • GPT-3.5-turbo with Persian-Sentence-Embedding-V3 achieves the highest context relevancy score (0.223), demonstrating its relative strength in capturing and incorporating contextual nuances, despite lower performance in other metrics.
    • Dorna with Dorna Embeddings also performs well (0.216), reflecting its capacity to balance context incorporation with other aspects of performance.
Observations and Insights:
  • Trade-offs between Models and Embeddings:
    • GPT-3.5-turbo with OpenAI Embeddings delivers the best performance in faithfulness, reflecting the general applicability and robustness of these embeddings across various datasets.
    • Conversely, Dorna with Dorna Embeddings exhibits superior performance in answer relevancy, emphasizing the importance of leveraging embeddings specifically designed for Persian text in achieving domain-specific objectives.
  • Performance Variability of Persian-Specific Embeddings: While Persian-Sentence-Embedding-V3 demonstrates strengths in context relevancy, its relatively lower performance in faithfulness and answer relevancy indicates the need for further optimization and training on diverse Persian datasets to improve its applicability for information retrieval tasks.
@@ VI-B Quality of Outputs
To assess the acceptability of the pipeline outputs from a human perspective, we conducted a qualitative evaluation involving 10 reviewers, including members of the University of Isfahan Artificial Intelligence Community and university students. The evaluators rated the generated answers based on clarity, coherence, and overall satisfaction. Feedback indicated that the majority of the answers were clear, well-structured, and effectively addressed the questions, demonstrating high acceptability.
Moreover, the consistency of the answers was noted, with evaluators observing a uniform level of quality across different questions. This consistency highlights the robustness of our pipeline, confirming its ability to produce reliable and high-quality answers suitable for practical use in a university setting.